# Darkonia at Colliders

### v1

@@ 1 Introduction
The existence of darkonia has significant effects on the dark matter phenomenology. Bound-state formation can enhance the dark matter annihilation cross section via the so-called Sommerfeld effect, with dramatic consequences on thermal freeze-out in the early universe and indirect dark matter detection today Hisano:2004ds ; Cirelli:2007xd ; Arkani-Hamed:2008hhe . In direct detection experiments, bound states of a dark matter particle and a nucleus can mediate resonant scattering, leading to an enhanced modulation signal Pospelov:2008qx ; Bai:2009cd . In this work, we focus on the impact of bound-state formation on the third main method of dark matter searches: production and decay at colliders. Unlike conventional dark matter searches for invisible particles through missing energy, the production of dark bound states at colliders typically results in a characteristic pattern of visible particles in the final state. Darkonium decays generate new signatures, many of which are not covered by current searches.
@@
Our analysis generalizes the results of a long history of studying bound states of hidden-sector particles. In supersymmetric theories, nearly mass-degenerate superpartners can form bound states whose decays produce characteristic signatures at high-energy colliders like the LHC Nappi:1981ft ; Barger:1988sp ; Martin:2008sv . Dark sectors with strongly interacting particles through a confining force similar to QCD can produce “dark showers” with signatures like emerging jets Schwaller:2015gea or semi-visible jets Cohen:2015toa . More recently, bound-state formation from a non-confining feeble force has been shown to predict interesting collider signatures An:2015pva ; Tsai:2015ugz ; Krovi:2018fdr . The models we consider share features with this last class of bound states with feeble interactions.
@@ 2.1 Dark sector models
$\mathcal{L}_{\rm portal}$ $=$ $-\mu_{S}\,S\,H^{\dagger}H-\lambda_{S}\,S^{2}\,H^{\dagger}H~{}+\mu^{2}H^{\dagger}H-\lambda(H^{\dagger}H)^{2}\quad(\mbox{F}_{\rm S}\;\mbox{model})$ (5)
@@
with $H$ the electroweak Higgs doublet in the Standard Model. For later purposes, we have included the Higgs potential in (5). Following our symmetry assumptions, we have neglected the possible parity-violating interaction $P\,H^{\dagger}H$.
@@
$\tan(2\,\xi)=\frac{2\,\eta\,s_{w}}{1-(\eta\,s_{w})^{2}-\delta}\ ,\qquad\delta=(1-\hat{\epsilon}^{2})\frac{m_{d}^{2}}{m_{Z}^{2}}\ ,\qquad m_{Z}=\frac{g\,v}{2\,c_{w}}\ ,$ (9)
with the Higgs vacuum expectation value $v=246$ GeV. In the limit of small kinetic mixing $\eta$ and mass ratio $\delta$, the mixing angle is approximated by $\xi\approx\epsilon\tan\theta_{w}$. The masses of the dark photon and $Z$ boson receive only second-order corrections in $\eta$ and $\delta$; we will neglect them. The photon remains massless. In this limit, the interaction eigenstates are related to the mass eigenstates by
@@ 2.2 Bound-state masses and wave functions
For the darkonium phenomenology at colliders, we consider bound states that are made of one dark fermion and one dark antifermion. Stable bound states can also be formed from two dark fermions or two dark antifermions if the binding force is scalar. Such bound states could have non-trivial cosmological consequences, but are more challenging to be produced at colliders. We do not consider them in this work. Furthermore, we concentrate on the lowest-lying bound states with zero orbital angular momentum $L=0$ (with one exception of including $L=1$ in the $\mbox{F}_{\rm S}$ model).
@@ 5 Conclusions and outlook
At future colliders, the scope of darkonium searches can be significantly extended. High-luminosity experiments like the FCC-ee FCC:2018evy or a muon collider Accettura:2023ked will allow to probe extremely rare signatures. For the scenarios considered in this work, these experiments will have an unprecedented sensitivity to darkonia with tiny couplings to the Standard Model. As a second benefit, future colliders will offer new search options for darkonia above the GeV scale. The FCC-ee or CEPC CEPCStudyGroup:2018ghi will extend the discovery potential of Belle II to much higher bound-state masses, providing an excellent environment to reconstruct signatures with displaced vertices and missing energy. At a muon collider, darkonia with scalar mediators can be produced through the new channels known for the SM Higgs boson, while darkonia with vector mediators benefit from efficient new production processes like vector boson fusion.

### v2

@@ 1 Introduction
The existence of darkonia has significant effects on the dark matter phenomenology. Bound-state formation (in addition to the Sommerfeld effect) can enhance the dark matter annihilation cross section vonHarling:2014kha , with dramatic consequences on thermal freeze-out in the early universe and indirect dark matter detection today Hisano:2004ds ; Cirelli:2007xd ; Arkani-Hamed:2008hhe ; Binder:2023ckj . In direct detection experiments, bound states of a dark matter particle and a nucleus can mediate resonant scattering, leading to an enhanced modulation signal Pospelov:2008qx ; Bai:2009cd . In this work, we focus on the impact of bound-state formation on the third main method of dark matter searches: production and decay at colliders. Unlike conventional dark matter searches for invisible particles through missing energy, the production of dark bound states at colliders typically results in a characteristic pattern of visible particles in the final state. Darkonium decays generate new signatures, many of which are not covered by current searches.
@@
Our analysis generalizes the results of a long history of studying bound states of hidden-sector particles. In supersymmetric theories, nearly mass-degenerate superpartners can form bound states whose decays produce characteristic signatures at high-energy colliders like the LHC Nappi:1981ft ; Barger:1988sp ; Martin:2008sv . Dark sectors with strongly interacting particles through a confining force similar to QCD can produce “dark showers” with signatures like emerging jets Schwaller:2015gea or semi-visible jets Cohen:2015toa . More recently, bound-state formation from a non-confining feeble force has been shown to predict interesting collider signatures Shepherd:2009sa ; An:2015pva ; Tsai:2015ugz ; Krovi:2018fdr . The models we consider share features with this last class of bound states with feeble interactions.
@@ 2.1 Dark sector models
$\mathcal{L}_{\rm portal}$ $=$ $-\mu_{S}\,S\,(H^{\dagger}H-\frac{v^{2}}{2})-\lambda_{S}\,S^{2}\,H^{\dagger}H~{}+\mu^{2}H^{\dagger}H-\lambda(H^{\dagger}H)^{2}\quad(\mbox{F}_{\rm S}\;\mbox{model})$ (5)
@@
where $H$ is the electroweak Higgs doublet in the Standard Model with vacuum expectation value (vev) $v=246$ GeV. For later purposes, we have included the Higgs potential in (5).
In (5), we have made a few assumptions and simplifications to the most general possible Lagrangian. We assume no vev for the scalar field $S$ and subtract the Higgs vev contribution from the first term, which would introduce a scalar vev through the linear term in $S$. The so-obtained interactions between the scalar $S$ and the Higgs boson $h$ are effectively the same as in a model where the dark scalar field obtains a vev, see e.g. Freitas:2015hsa . We neglect a possible linear term in $S$ and self-interaction terms $S^{3}$ and $S^{4}$. [^1] In Sec. 3.1, we comment on the impact of self-interactions on bound-state decays. These assumptions are purely phenomenologically motivated. They do not impact the dominant darkonium production channels at colliders, while allowing us to study the darkonium phenomenology in terms of a few parameters.
Following our symmetry assumptions, we have also neglected the possible parity-violating interaction $P\,H^{\dagger}H$ in (6).
@@
$\tan(2\,\xi)=\frac{2\,\eta\,s_{w}}{1-(\eta\,s_{w})^{2}-\delta}\ ,\qquad\delta=(1-\hat{\epsilon}^{2})\frac{m_{d}^{2}}{m_{Z}^{2}}\ ,\qquad m_{Z}=\frac{g\,v}{2\,c_{w}}\ .$ (9)
In the limit of small kinetic mixing $\eta$ and mass ratio $\delta$, the mixing angle is approximated by $\xi\approx\epsilon\tan\theta_{w}$. The masses of the dark photon and $Z$ boson receive only second-order corrections in $\eta$ and $\delta$; we will neglect them. The photon remains massless. In this limit, the interaction eigenstates are related to the mass eigenstates by
@@ 2.2 Bound-state masses and wave functions
For the darkonium phenomenology at colliders, we consider bound states that are made of one dark fermion and one dark antifermion. Stable bound states can also be formed from two dark fermions or two dark antifermions if the binding force is scalar. Such bound states could have non-trivial cosmological consequences, but are more challenging to be produced at colliders. The reason is that, to conserve the dark fermion number, a total of four particles must be produced — two dark fermions and two dark antifermions. Since four-particle production at colliders is phase-space suppressed, fermion-fermion or antifermion-antifermion bound states are less likely to be produced.
@@
2.3 Dark matter relic abundance
Throughout this work, we will be agnostic about the relic abundance of both the dark matter constituents and the dark matter bound states. In this way, our predictions of darkonium production and decay can be generally applied to scenarios with or without a dark matter context. In this section we comment on freeze-out scenarios that could explain the dark matter relic abundance through early-universe dynamics.
In general, the dark matter abundance today can be a mixture of dark fermion constituents and (stable) bound states. In the ${\rm F}_{\rm S}$ and ${\rm F}_{\rm P}$ models, the lightest $P_{d}$- and $C_{d}$-odd states are stable and contribute to the total dark matter energy density.
The thermal freeze-out of the dark constituents follows the standard calculation Gondolo:1990dk , including potential effects of bound state formation vonHarling:2014kha . In the absence of bound state effects, the leading annihilation channel of a dark fermion-antifermion pair is into two force mediators. Using the ${\rm F}_{\rm S}$ model as an example, the annihilation rate for $m_{\chi}\gg m_{S}$ reads
$\sigma(\chi\bar{\chi}\rightarrow SS)\,v=\frac{3g_{d}^{2}}{128\pi m_{\chi}^{2}}\,v^{2}$ (20)
and is $p$-wave suppressed. Here, $v\sim 0.3$ c is the dark fermion speed at around the freeze-out temperature. To satisfy the total relic abundance, one has $3g_{d}^{2}/(128\pi m_{\chi}^{2})\approx 5.5$ pb Bai:2012qy or $m_{\chi}/g_{d}\approx 735\,$GeV. The formation of bound states can significantly enhance the annihilation cross section and allow for much lighter dark matter candidates. Lighter dark matter is also favored in the scenario where the constituents only make up a fraction of the total dark matter abundance.
The abundance of stable bound states can be calculated in a similar way as for hydrogen formation. In analogy with the recombination process $e^{-}p\rightarrow H\gamma$, the relic bound state abundance is determined by processes like $\chi\bar{\chi}\rightarrow\eta_{d}S$ in the ${\rm F}_{\rm S}$ model. Since the mediator $S$ is massive, the rate of bound state formation depends on the model parameters, e.g. the relation between binding energy and $S$-particle mass. The formation of fermion-fermion bound states could further complicate the total relic abundance calculation. Given the scope of this work, we do not explore the origin of the bound state abundance in detail.
Finally, the LHC phenomenology places greater emphasis on the portal coupling, which is not as crucial for bound state formation in the early universe. Therefore, one has the flexibility to reduce the coupling to comply with current LHC constraints, while still expecting detectable signals at the HL-LHC and future experiments.
@@ 5 Conclusions and outlook
At future colliders, the scope of darkonium searches can be significantly extended. High-luminosity experiments like the FCC-ee FCC:2018evy or a muon collider Accettura:2023ked will allow to probe extremely rare signatures. For the scenarios considered in this work, these experiments will have an unprecedented sensitivity to darkonia with tiny couplings to the Standard Model. As a second benefit, future colliders will offer new search options for darkonia well above the weak scale Bottaro:2021srh . The FCC-ee or CEPC CEPCStudyGroup:2018ghi will extend the discovery potential of Belle II to much higher bound-state masses, providing an excellent environment to reconstruct signatures with displaced vertices and missing energy. At a muon collider, darkonia with scalar mediators can be produced through the new channels known for the SM Higgs boson, while darkonia with vector mediators benefit from efficient new production processes like vector boson fusion.